given that a=[
1 0 3
2 1 -1
1 -1 1 ],compute the value of (a^6-5a^5+8a^4-2a^3-9a^2+31a-36 I ) using cayley-hamilton theorem
1
Expert's answer
2020-12-20T18:36:08-0500

Characteristics equation is "|A-\lambda I|=0".

"\begin{vmatrix}
 1&0&3 \\
 2&1&-1\\
 1&-1&1\\
\end{vmatrix}=0"

"(1-\lambda) [(1-\lambda) ^2-1]+3[-2-(1-\lambda) ]=0\\
(1-\lambda) ^3-1+\lambda+3(-3+\lambda) =0\\
(1-\lambda) ^3-1+\lambda-9+3\lambda=0\\
(1-\lambda)^3+4\lambda-10=0\\
1-3\lambda+3\lambda^2-\lambda^3+4\lambda-10=0\\
-\lambda^3+3\lambda^2+\lambda-9=0\\
\lambda^3-3\lambda^2-\lambda+9=0"

By Cayley-Hamilton Theorem, "A^3-3A^2-A+9=0."

"\implies A^3-3A^2=A-9I."


"A^6-5A^5+8A^4-2A^3-9A^2+31A-31I\\
=A^6-3A^5-2A^5+6A^4+2A^4-6A^3+4A^3-12A^2+3A^2+31A-36I\\
=A^3(A^3-3A^2)-2A^2(A^3-3A^2)+2A(A^3-3A^2)+4(A^3-3A^2)+3A^2+31A-36I\\
=(A^3-3A^2)(A^3-2A^2+2A+4I)+3A^2+31A-36I\\
=(A^3-3A^2)(A^3-3A^2+A^2-A+3A+9I-5I)+3A^2+31A-36I\\
=(A^3-3A^2)(A^3-3A^2-A+9I+A^2+3A-5I)+3A^2+31A-36I\\"

But, "(A^3-3A^2)=A-9I, A^3-3A^2-A+9I=0"

"=(A-9I)(0+A^2+3A-5)+3A^2+31A-36I\\
=A^3+3A^2-5A-9A^2-27A+45I+3A^2+31A-36I\\
=A^3-3A^2-A+9I\\
=0"